# Leveraging Large Language Model for Information Retrieval-based Bug Localization

### v1

@@ IV Experiement Setup
This section presents the implementation details, research questions, baseline techniques, benchmark dataset, and evaluation metrics.
IV-A Implementation
@@
Similar to most of the existing bug localization techniques [20, 5, 21], GenLoc is implemented in Java since it is one of the most popular programming languages. However, the approach is easily extendable to other languages, as it leverages the Tree-sitter library [36] for source code parsing, which supports more than ten programming languages. To compute semantic similarity between bug reports and source code files, GenLoc employs OpenAI’s text-embedding-3-small model due to its cost-effectiveness (\$0.02 per 1M tokens) [37]. Following established methods [38, 39], the source code is divided into chunks of 300 tokens. For each bug report, the top 50 source files are retrieved based on their embedding similarity to the report.
@@
To improve efficiency and minimize embedding costs, embeddings for all source files are generated and stored in ChromaDB [40] when processing the first bug in each project. Since different bugs occur in different versions, for each subsequent bug, only the embeddings of files that have been added, modified, deleted, or renamed across versions are updated. For inference, GPT-4o mini by OpenAI is used due to its cost-effectiveness (\$0.15 per 1M input tokens and \$0.60 per 1M output tokens) [41]. All model parameters are kept at their default settings to avoid the computational cost of hyperparameter tuning. In particular, the temperature parameter is left at its default value (1.0) (rather than being set to 0) to enable diverse reasoning paths during inference [42, 43]. A temperature of 0 significantly reduces variability by consistently selecting the highest-probability tokens, which can lead the model to overlook less obvious but relevant files. Our preliminary analysis also showed that the performance remained mostly unchanged regardless of the temperature setting. This aligns with prior findings that changes in temperature between 0.0 and 1.0 do not have a significant effect on LLM’s performance for problem-solving tasks [44]. To account for LLM non-determinism, the experiments were executed three times, and the average results were reported, as followed in [45]. The experiments were conducted on a MacBook Pro with an M1 Max chip, 32GB RAM, running macOS Sonoma 14.3.1.
IV-B Research Questions
@@
RQ1: How does GenLoc compare to other bug localization techniques?
This question analyzes the effectiveness of GenLoc by comparing its performance with existing state-of-the-art techniques using standard evaluation metrics.
RQ2: How do different components of GenLoc contribute to the overall performance?
This question examines the impact of individual components-such as the embedding-based retriever and the LLM reasoning module-on the final bug localization accuracy.
IV-C Baseline, Benchmark Dataset and Evaluation Metrics
@@
GenLoc is compared against a set of techniques with publicly available and workable implementations. We selected five widely studied IRBL techniques that span both traditional and deep learning-based methods: VSM, BugLocator, BLUiR, BRTracer, and DreamLoc. VSM computes cosine similarity between the bug report and source code based on term weights derived using Term Frequency–Inverse Document Frequency (TF-IDF) [46]. BugLocator extends VSM by incorporating file length normalization and historical bug-fix frequency to prioritize more fault-prone files [9]. BLUiR assigns different weights to structural program elements such as class and method names, and computes their similarity to the bug report content separately [47]. BRTracer divides source files into smaller segments (chunks) and integrates additional signals such as stack traces and historical fix data to rank files [29]. DreamLoc leverages a Wide and Deep architecture. The Wide component captures software-specific statistical features (e.g., fix frequency and recency), while the Deep component applies a relevance model for fine-grained similarity scoring [5]. For DreamLoc, the original implementation is used. For BugLocator, BLUiR, and BRTracer, the implementations provided in [1] and for VSM the implementation provided in [46] are adopted since the original implementations are not publicly available. While we aimed to incorporate several other IRBL techniques in the evaluation, they could not be included due to various limitations. For instance, DNNLOC [16] and DeepLoc [17] do not have a publicly released implementation. Although RLocator [21] provides a replication package, it is incomplete due to missing source code. AdaptiveBL [48] could not be executed because of unresolved dependency issues, which also affects FLIM [19], as it builds on AdaptiveBL.
@@
For comparison, the dataset by Ye et al. [20] is used, since it is the most widely used dataset in prior work [5, 13, 48]. It contains over 22,000 bugs from six large, open-source projects, as shown in Table I. Following prior papers [49, 21], 60% data is used as historical (training) data and the most recent 40%—consisting of 9,097 bugs—is used for evaluation.
TABLE I: Dataset Description
Project Time Range # Bug Reports # Java Files LOC
@@
AspectJ Mar 2002 - Jan 2014 593 6,879 4,439 2,076 699,250 515,153 216,387
Birt Jun 2005 - Dec 2013 4,178 9,697 6,841 1,700 2,322,074 1,549,525 412,795
Eclipse Oct 2001 - Jan 2014 6,495 6,243 3,454 382 1,156,041 637,158 47,348
JDT Oct 2001 - Jan 2014 6,274 10,544 8,184 2,294 922,812 609,378 133,323
SWT Feb 2002 - Jan 2014 4,151 2,795 2,056 1,037 934,357 639,216 301,698
Tomcat Jul 2002 - Jan 2014 1,056 2,042 1,552 924 487,701 413,472 270,046
Similar to other work [9, 50], the following three metrics are used for evaluation:
1. Accuracy@k: It denotes the number of bugs for which at least one of the actual buggy files is ranked within the top-k positions (k = 1, 5, 10) of the ranked list. A higher accuracy@k value indicates a better performance of the bug localization technique. Similar to other studies [5, 19], the value of k is set to 1, 5 and 10.
@@
$MRR@k=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}\frac{1}{\text{rank}_{i}}&\text{if }\text{rank}_{i}\leq k\\ 0&\text{otherwise}\end{cases}$ (1)
where, $N$ denotes the total number of bug reports. $\text{rank}_{i}$ refers to the rank of the first correct buggy file for the i-th bug report among the top-k results. If no actual buggy file appears within the top-k results for a bug report, the reciprocal rank is defined as 0. The higher the value of MRR@k, the better the technique. Following existing studies [50, 51], the value of k is set to 10. An MRR@10 value of 0.2 indicates that, on average, the first relevant file is ranked around position 5 within the top 10 results.
@@
$AP_{i}=\frac{\sum_{j=1}^{M}P(j)\times\text{Rel}(j)}{\text{Number of relevant source files}}$ (2)
where, $M$ is the number of retrieved files in the top-k ranked list for the i-th bug report (i.e., $M$=$k$). $Rel(j)$ denotes whether the file at position $j$ is relevant or not. $P(j)$ indicates the precision at rank $j$ and computed as follows:
$P(j)=\frac{\text{Number of relevant files in top-}j\text{ positions}}{j}$ (3)
@@
$MAP=\frac{\sum_{i=1}^{N}AP_{i}}{N}$ (4)
where, $N$ denotes the total number of bug reports. A higher MAP@k value indicates better ranking quality of the buggy files. Following prior studies [50, 51], the value of k is set to 10. For example, a MAP@10 value of 0.3 implies that, on average, 30% of the actual buggy files are retrieved with high precision within the top 10 results.
V Result Analysis
@@
V-A RQ1: Overall Performance of GenLoc
GenLoc is evaluated against five state-of-the-art baseline techniques (DreamLoc, BRTracer, BLUiR, BugLocator, and VSM) across six widely-used benchmark projects (AspectJ, Birt, Eclipse, JDT, SWT, and Tomcat). Tables II and III present the overall and project-wise results, respectively.
TABLE II: Overall Performance of Bug Localization Techniques
@@ V-A RQ1: Overall Performance of GenLoc
VSM 4.01 9.78 14.22 0.05 0.07
V-A1 Overall Results
Table II shows that GenLoc demonstrates strong overall performance across all evaluation metrics compared to prior bug localization techniques. On average, GenLoc achieves the highest Accuracy@1 (43.19%), substantially outperforming the second-best technique (BRTracer with 26.84%) by a relative margin of over 60%. This indicates that GenLoc can correctly identify the buggy file at the top position in a significant portion of bug reports. Similarly, GenLoc leads in Accuracy@5 (62.35%) and Accuracy@10 (67.87%), further affirming its ability to prioritize buggy files effectively within the top candidates. In terms of ranking quality, GenLoc achieves the highest scores in both MAP@10 (0.41) and MRR@10 (0.51), outperforming the next-best MAP@10 (DreamLoc with 0.30) and MRR@10 (BRTracer and DreamLoc with 0.37) by 36.67% and 37.84%, respectively, confirming its superiority in ranking relevant files near the top of the result list.
@@ V-A1 Overall Results
TABLE III: Project-wise Performance Comparison of Bug Localization Techniques
@@
VSM 9.24 23.70 35.55 0.13 0.16
V-A2 Project-wise Results
Table III shows that GenLoc achieves the highest Accuracy@1 across all six projects. Notably, it demonstrates substantial improvements—ranging from 30% to 80%—over the best-performing baseline on projects such as AspectJ, Eclipse, JDT, SWT, and Tomcat. At Accuracy@5, GenLoc continues to outperform all baselines in five out of six projects, most notably achieving 83.73% on Tomcat, 66.72% on JDT, and 65.99% on Eclipse. A similar trend is observed for Accuracy@10, where GenLoc continues to outperform or match other techniques.
Regarding ranking quality, GenLoc achieves the highest MAP@10 and MRR@10 scores on five of the six projects. For instance, on AspectJ, GenLoc achieves an MRR of 0.53, far exceeding that of the second-best technique (BRTracer with 0.38). Similar trends are observed in the MAP scores on projects such as Eclipse, JDT, SWT, and Tomcat, where GenLoc consistently outperforms all baselines, further demonstrating its strength in producing effective ranked outputs.
@@ V-A2 Project-wise Results
Figure 3: Bug Report from Birt Project.
Although GenLoc achieves the best Accuracy@1 (19.71%) on Birt project, its Accuracy@5 and Accuracy@10 are lower compared to DreamLoc. Additionally, GenLoc attains the second-highest MAP (0.20) and MRR (0.26), with DreamLoc slightly outperforming it on both metrics. To investigate this performance gap, the authors manually analyzed several Birt bug reports that GenLoc failed to localize and found a common pattern that many of these reports contained “steps to reproduce” information. While such descriptions are helpful for developers [52, 53], they primarily focus on user actions (e.g., clicking buttons or selecting options), which rarely align directly with the source code, as shown in Fig. 3. This disconnection makes it difficult for GenLoc to extract meaningful semantic clues from the report and map them to relevant files.
To systematically verify this observation, reproduction related keywords are extracted from all bug reports following prior work [53]. As shown in Table IV, 63.85% of Birt’s bug reports include such reproduction-related content, which is substantially higher than any other project (e.g., 3.80% in AspectJ, 2.62% in Eclipse, and 2.91% in JDT). This discrepancy contributes to the semantic mismatch between bug reports and source files, which hampers GenLoc’s ability to accurately rank relevant files. Furthermore, a chi-square test revealed a statistically significant relationship between the presence of reproduction information and the failure to localize bugs in the Birt project across all three GenLoc trials (p-value = 0). In contrast, DreamLoc considers additional features such as bug-fixing history of a file, cyclomatic complexity, apart from similarity matching. These supplementary features helped DreamLoc outperform GenLoc on Birt. However, when DreamLoc is executed without these additional features, its performance also drops significantly, as shown in Table LABEL:dreamloc.
TABLE IV: Distribution of bug reports with reproduction-related keywords
Project Total Bugs
Reproduction-related
Keywords
Percentage (%)
@@
TABLE V: Comparison of GenLoc and DreamLoc (without additional features) on the Birt Project
@@
V-A3 Overlap Analysis
@@
Apart from Accuracy@k, MRR@10 and MAP@10, the number of unique and overlapping bugs localized by each approach are examined. In this analysis, a bug is considered successfully localized if the correct buggy file appears within the top 10 ranked results (i.e., Accuracy@10), aligning with a prior study showing that over 98% of developers consider localization results successful when buggy elements are ranked within the top ten positions [54]. Table VI presents the number of total and unique bugs localized by each technique under this criterion. Since GenLoc is a non-deterministic approach, its output varies across different executions due to the inherent randomness in language model sampling. To fairly evaluate GenLoc’s overall capability and compare it with existing deterministic baselines, the union of all bugs localized across multiple GenLoc trials is considered. This union represents the total set of unique bugs that GenLoc is able to localize in at least one run, capturing the full potential of the approach rather than the outcome of a single arbitrary execution.
TABLE VI: Result of Overlap Analysis
Technique Localized Bugs Overlapping Bugs Unique Bugs
GenLoc 6824 6234 590
DreamLoc 5889 5561 328
BRTracer 5283 5250 33
BLUiR 5286 5200 86
BugLocator 5055 5038 17
VSM 1107 1096 11
@@
Among all approaches, GenLoc exhibits the highest overall coverage with 6,824 localized bugs, of which 590 are uniquely identified by GenLoc. This indicates that GenLoc not only localizes more bugs overall but also captures a unique subset of bugs missed by all other methods, highlighting its potential to complement existing methods. This is likely due to its ability to reason over natural language and semantic context, which prior techniques lack.
Bug ID: 55125
Summary: Tomcat does not shut down if LifecycleListener throws an exception
Description: Created attachment 30467 Logs I needed to shut down Tomcat if an application fails to start. I was advised that a LifecycleListener could shut down Tomcat by throwing an IllegalStateException from its lifecycleEvent method. However, this produces severe errors in the logs, but Tomcat does not shut down. I saw this behaviour on Tomcat 7.0.41 and 7.0.40 using the test web application from here: https://github.com/cloudfoundry/java-test-applications and v1.0.0 of the listener from here: https://github.com/cloudfoundry/java-buildpack-support (v1.1.0 of the listener was re-coded to halt the JVM). I attach catalina.out. Mark Thomas suggested I raise this bug, so he believes that Tomcat should shut down.
Figure 4: Bug Report from Tomcat Project.
V-A4 Case Study
A sample bug [^2] https://bz.apache.org/bugzilla/show_bug.cgi?id=55125 from Tomcat project is shown in Fig. 4, where all prior techniques such as BugLocator, BLUiR, and DreamLoc failed to localize the bug due to vocabulary gap between the bug report and source code. In contrast, GenLoc could place the correct buggy file Catalina.java at the top (rank 1). It first understood that the core issue in the bug report was related to Tomcat failing to shut down when a LifecycleListener throws an exception. Based on this understanding, it identified files related to lifecycle events by invoking get_candidate_filenames() and search_file(). It then used get_method_signatures_of_a_file() to inspect the structure of these files and locate shutdown-related methods. Finally, by analyzing the method bodies using get_method_body(), it examined functions like start() and stop() in Catalina.java, ultimately pinpointing it as the buggy file. This example highlights GenLoc’s strength in reasoning about semantic intent and program behavior, going beyond surface-level text matching.
V-B RQ2: Ablation Study
To better understand the contribution of different components within GenLoc’s bug localization pipeline, an ablation study is conducted. In particular, the goal is to examine how embedding-based file retrieval and LLM-based iterative reasoning interact to improve bug localization performance. By selectively disabling or isolating these components, their individual and combined impact are examined. For this purpose, the following three configurations are used:
1. GenLoc: This is the complete version of the proposed approach, where the LLM is provided with five functions, including get_candidate_filenames() to utilize embedding-recommended files.
2. Embedding-Only: This variant evaluates the standalone performance of the embedding-based retrieval system, where source files are ranked solely based on their semantic similarity to the bug report, without any reasoning or iteration by the LLM.
3. GenLoc-NoEmbed: In this configuration, the LLM operates without access to the embedding-based recommended files, i.e., the get_candidate_filenames() function is removed. The model must infer filenames directly from the bug report and continue reasoning with the remaining four functions.
Table VII presents the average results across all projects. GenLoc achieves the best performance on all metrics, demonstrating the strength of combining embedding-based retrieval with iterative LLM analysis. The Embedding-Only variant obtains only 32.59% at Accuracy@1, indicating that retrieval alone is insufficient without deeper reasoning. A sample case [^3] https://bugs.eclipse.org/bugs/show_bug.cgi?id=232757 is shown in Fig. 5, where Embedding-Only was unable to localize the correct file, while GenLoc succeeded. Due to space constraints, a selection of representative lines from the bug report is presented. The Embedding-Only approach failed to localize the correct buggy file due to excessive noise in the bug report, which mentioned over 50 unique filenames, most of which were unrelated to the fault. GenLoc overcame this challenge by first identifying key method names from the stack trace (e.g., zoomOut() and render()), then searching for files containing these methods. It subsequently examined the method bodies to assess their relevance to the failure. This structured reasoning enabled GenLoc to correctly localize the buggy file, AutoScale.java, despite the presence of substantial noise in the report.
TABLE VII: Result of Ablation Study

### v2

@@ 4. Experiement Setup
This section presents the implementation details, research questions, baseline techniques, benchmark dataset and evaluation metrics.
4.1. Implementation
@@
Similar to most of the existing techniques (Ye et al., 2015; Qi et al., 2021; Chakraborty et al., 2024), GenLoc focuses on Java projects since it is one of the most popular programming languages. However, the approach is easily extendable to other languages, as it leverages the Tree-sitter library (tre, [n. d.]) for source code parsing, which supports more than ten programming languages. To compute semantic similarity between bug reports and source code files, GenLoc employs OpenAI’s text-embedding-3-small model due to its cost-effectiveness (\$0.02 per 1M tokens) (tex, [n. d.]). Following Amazon Bedrock’s recommendation (aws, [n. d.]), we use a chunk size of 300 tokens to balance between granularity and context—capturing relevant code segments while preserving surrounding information for meaningful semantic comparison. For each bug report, the top 50 source files are retrieved based on their embedding similarity to the report (Li et al., 2025a).
@@
To improve efficiency and minimize embedding costs, embeddings for all source files are generated and stored in ChromaDB (chr, [n. d.]) when processing the first bug in each project. Since different bugs occur in different versions, for each subsequent bug, only the embeddings of files that have been added, modified, deleted, or renamed across versions are updated in the ChromaDB. For inference, GPT-4o mini by OpenAI is used due to its cost-effectiveness (\$0.15 per 1M input tokens and \$0.60 per 1M output tokens) (gpt, [n. d.]). All model parameters are kept at their default settings to avoid the computational cost of hyperparameter tuning. In particular, the temperature parameter is left at its default value (1.0) (rather than being set to 0) to enable diverse reasoning paths during inference (Kang et al., 2023; Zhu et al., 2024). A temperature of 0 significantly reduces variability by consistently selecting the highest-probability tokens, which can lead the model to overlook less obvious but relevant files. Our preliminary analysis also showed that the performance remained mostly unchanged regardless of the temperature setting. This aligns with prior findings that changes in temperature between 0.0 and 1.0 do not have a significant effect on LLM’s performance for problem-solving tasks (Renze, 2024). To account for LLM non-determinism, the experiments were executed three times and the average results were reported, as followed in (Zhang et al., 2024).
4.2. Research Questions
@@
  • RQ1: How does GenLoc perform compared to traditional and deep learning-based IRBL techniques?
@@
  • RQ2: How does GenLoc perform compared to other LLM-based IRBL techniques?
  • RQ3: How do different components of GenLoc contribute to the overall performance?
  • RQ4: How well does GenLoc perform on previously unseen bugs?
4.3. Baseline Techniques
To provide a comprehensive evaluation of GenLoc, we compare it against a diverse set of file-level IRBL techniques with publicly available and executable implementations. The baselines cover both traditional, deep learning-based methods and recent LLM-based approaches.
@@ 4.3. Baseline Techniques
We selected four widely studied IRBL techniques from the traditional and deep learning categories: BugLocator (Zhou et al., 2012), BLUiR (Saha et al., 2013), BRTracer (Wong et al., 2014) and DreamLoc (Qi et al., 2021). BugLocator uses revised Vector Space Model (rVSM) by incorporating file length normalization and historical bug-fix frequency to prioritize more fault-prone files (Zhou et al., 2012). BLUiR assigns different weights to structural program elements such as class and method names, and computes their similarity to the bug report content separately (Saha et al., 2013). BRTracer divides source files into smaller segments (chunks) and integrates additional signals such as stack traces and historical fix data to rank files (Wong et al., 2014). DreamLoc leverages a Wide and Deep architecture. The Wide component captures software-specific statistical features (e.g., fix frequency and recency), while the Deep component applies a relevance model for fine-grained similarity scoring (Qi et al., 2021).
For DreamLoc, the original implementation is used. For BugLocator, BLUiR and BRTracer, the implementations provided in (Lee et al., 2018) are adopted, since the original implementations are unavailable. While we aimed to incorporate several other IRBL techniques in the evaluation, they could not be included due to implementation-related issues. For instance, DNNLOC (Lam et al., 2017) and DeepLoc (Xiao et al., 2019) do not have a publicly released implementation. Although RLocator (Chakraborty et al., 2024) provides a replication package, it is incomplete due to missing source code. AdaptiveBL (Fejzer et al., 2021) and its successor FLIM (Liang et al., 2022) could not be executed due to unresolved dependency issues.
In addition, we compare GenLoc against three LLM-based approaches: BRaIn (Samir and Rahman, 2025), LLM-BL (Li et al., 2025a) and Agentless (Xia et al., 2025). BRaIn retrieves top-K files using lexical search, identifies relevant methods from those files using an LLM and refines the results through keyword-based query expansion and reranking (Samir and Rahman, 2025). LLM-BL expands the bug report using an LLM, retrieves candidate files through hybrid (lexical and semantic) search and presents the shortlisted files along with their relevant code lines to the LLM for reranking (Li et al., 2025a). Agentless is the leading open-source issue-fixing technique and has been adopted by OpenAI as the standard approach for showcasing real-world coding performance (Xia et al., 2025; Chang et al., 2025). It localizes bugs through three steps: (i) representing the code base as a tree-like structure and prompting the LLM to rank suspicious files, (ii) using embeddings to retrieve additional candidates, and (iii) combining both results into the final list of suspicious files. For both BRaIn and LLM-BL, we use their original implementations, while for Agentless, we extract its file-level bug localization component and adapt it (e.g., code parsing) for Java projects, since the original version only supports Python.
4.4. Benchmark Dataset
@@
To evaluate GenLoc, we employ two complementary benchmark datasets: the widely used Ye et al. dataset (Ye et al., 2015) and the more recent GHRB dataset (Lee et al., 2024). For RQ1, we use the Ye et al. dataset which is the most widely adopted benchmark in prior IRBL studies (Qi et al., 2021; Ye et al., 2016; Fejzer et al., 2021). This dataset includes over 22,000 bugs from six large, open-source projects (Table 1), which makes it suitable for training deep-learning based approach DreamLoc. Following prior work (Han et al., 2023a; Chakraborty et al., 2024), 60% data is used as historical (training) data and the most recent 40%—consisting of 9,097 bugs—is used for evaluation.
Table 1. Ye et al. Dataset Description
Project # of Bug Reports # of Java Files LOC
@@
AspectJ 593 6,879 4,439 2,076 699,250 515,153 216,387
Birt 4,178 9,697 6,841 1,700 2,322,074 1,549,525 412,795
Eclipse 6,495 6,243 3,454 382 1,156,041 637,158 47,348
JDT 6,274 10,544 8,184 2,294 922,812 609,378 133,323
SWT 4,151 2,795 2,056 1,037 934,357 639,216 301,698
Tomcat 1,056 2,042 1,552 924 487,701 413,472 270,046
For RQ2 and RQ3, we use the GHRB (GitHub Recent Bugs) dataset, which is specifically designed for evaluating LLM-based approaches (Lee et al., 2024). It contains 131 recent bugs from 16 projects (e.g., Apache Dubbo, OpenAPI Generator), as shown in Table 2. The size of the dataset enables us to execute LLM-based approaches three times to account for their non-deterministic behavior. To further validate our findings, we also experiment with the latest 300 bugs (50 from each project) from the Ye et al. dataset. For RQ4, we use the subset of GHRB bug reports submitted after the knowledge cut-off date of GPT-4o-mini (October 1, 2023) (gpt, [n. d.]), leaving 49 unseen bug reports for evaluation.
Table 2. GHRB Dataset Description
# of Projects # of Bug Reports # of Java Files LOC
Max Median Min Max Median Min
16 131 12,025 1,977 93 1,874,139 299,697 14,759
4.5. Evaluation Metrics
Similar to other work (Zhou et al., 2012; Rahman and Roy, 2018), the following three metrics are used for evaluation:
1. Accuracy@k: It denotes the number of bugs for which at least one of the actual buggy files is ranked within the top-k positions (k = 1, 5, 10) of the ranked list. A higher accuracy@k value indicates a better performance of the bug localization technique. Similar to other studies (Qi et al., 2021; Liang et al., 2022), the value of k is set to 1, 5 and 10.
@@ 4.5. Evaluation Metrics
(1) $MRR@k=\frac{1}{N}\sum_{i=1}^{N}\begin{cases}\frac{1}{\text{rank}_{i}}&\text{if }\text{rank}_{i}\leq k\\ 0&\text{otherwise}\end{cases}$
where, $N$ denotes the total number of bug reports. $\text{rank}_{i}$ refers to the rank of the first correct buggy file for the i-th bug report among the top-k results. If no actual buggy file appears within the top-k results for a bug report, the reciprocal rank is defined as 0. The higher the value of MRR@k, the better the technique. Following existing studies (Rahman and Roy, 2018; Kim and Lee, 2019), the value of k is set to 10. An MRR@10 value of 0.2 indicates that, on average, the first relevant file is ranked around position 5 within the top 10 results.
@@
(2) $AP_{i}=\frac{\sum_{j=1}^{M}P(j)\times\text{Rel}(j)}{\text{Number of relevant source files}}$
where, $M$ is the number of retrieved files in the top-k ranked list for the i-th bug report (i.e., MM=$k$). $Rel(j)$ denotes whether the file at position $j$ is relevant or not. $P(j)$ indicates the precision at rank $j$ and computed as follows:
(3) $P(j)=\frac{\text{Number of relevant files in top-}j\text{ positions}}{j}$
@@
(4) $MAP=\frac{\sum_{i=1}^{N}AP_{i}}{N}$
where, $N$ denotes the total number of bug reports. A higher MAP@k value indicates better ranking quality of the buggy files. Following prior studies (Rahman and Roy, 2018; Kim and Lee, 2019), the value of k is set to 10. For example, a MAP@10 value of 0.3 implies that, on average, 30% of the actual buggy files are retrieved with high precision within the top 10 results.
5. Result Analysis
@@
5.1. RQ1: Comparison with Traditional and Deep Learning-based IRBL Techniques
GenLoc is evaluated against four traditional and deep-learning based baseline techniques (DreamLoc, BRTracer, BLUiR and BugLocator) across six widely-used benchmark projects (AspectJ, Birt, Eclipse, JDT, SWT and Tomcat). Tables 3 and 4 present the overall and project-wise results, respectively.
Table 3. Comparison of GenLoc and Non-LLM-based Bug Localization Techniques
@@ 5.1. RQ1: Comparison with Traditional and Deep Learning-based IRBL Techniques
Table 3 shows that GenLoc demonstrates strong overall performance across all evaluation metrics compared to prior bug localization techniques. On average, GenLoc achieves the highest Accuracy@1 (44.01%), substantially outperforming the second-best technique (BRTracer with 26.84%) by a relative margin of over 63%. This indicates that GenLoc can correctly identify the buggy file at the top position in a significant portion of bug reports. Similarly, GenLoc leads in Accuracy@5 (63.27%) and Accuracy@10 (68.28%), further affirming its ability to prioritize buggy files effectively within the top candidates. In terms of ranking quality, GenLoc achieves the highest scores in both MAP@10 (0.41) and MRR@10 (0.52), outperforming the next-best MAP@10 (DreamLoc with 0.30) and MRR@10 (BRTracer and DreamLoc with 0.37) by 36.67% and 40.54%, respectively, confirming its superiority in ranking relevant files near the top of the result list.
@@
Table 4. Project-wise Performance Comparison of Bug Localization Techniques
@@
Table 4 shows that GenLoc achieves the highest Accuracy@1 across all six projects. Notably, it demonstrates substantial improvements–ranging from 36% to 91%–over the best-performing baseline on projects such as AspectJ, Eclipse, JDT, SWT and Tomcat. At Accuracy@5, GenLoc continues to outperform all baselines in five out of six projects, most notably achieving 83.89% on Tomcat, 67.48% on JDT, and 66.95% on Eclipse. A similar trend is observed for Accuracy@10, where GenLoc continues to outperform other techniques.
Regarding ranking quality, GenLoc achieves the highest MAP@10 and MRR@10 scores on five of the six projects. For instance, on AspectJ, GenLoc achieves an MRR of 0.55, far exceeding that of the second-best technique (BRTracer with 0.38). Similar trends are observed in the MAP scores on projects such as Eclipse, JDT, SWT and Tomcat, where GenLoc consistently outperforms all baselines, further demonstrating its strength in producing effective ranked outputs.
Although GenLoc achieves the best Accuracy@1 (19.67%) on Birt project, its Accuracy@5 and Accuracy@10 are lower compared to DreamLoc. Additionally, GenLoc attains the second-highest MAP (0.19) and MRR (0.26), with DreamLoc slightly outperforming it on both metrics. To investigate this performance gap, the authors manually analyzed several Birt bug reports that GenLoc failed to localize and found a common pattern that many of these reports contained “steps to reproduce” information. While such descriptions are helpful for developers (Bettenburg et al., 2008; Soltani et al., 2020), they primarily focus on user actions (e.g., clicking buttons or selecting options), which rarely align directly with the source code, as shown in Fig. 3. This disconnection makes it difficult for GenLoc to extract meaningful semantic clues from the report and map them to relevant files.
@@
Figure 3. Bug Report from Birt Project.
To systematically verify this observation, reproduction related keywords are extracted from all bug reports following prior work (Soltani et al., 2020). As shown in Table 5, 63.85% of Birt’s bug reports include such reproduction-related content, which is substantially higher than any other project (e.g., 3.80% in AspectJ, 2.62% in Eclipse, and 2.91% in JDT). This discrepancy contributes to the semantic mismatch between bug reports and source files, which hampers GenLoc’s ability to accurately rank relevant files. Furthermore, a chi-square test revealed a statistically significant relationship between the presence of reproduction information and the failure to localize bugs in the Birt project across all three GenLoc trials (p-value = 0). In contrast, DreamLoc considers additional features such as bug-fixing history of a file, cyclomatic complexity, apart from similarity matching. These supplementary features helped DreamLoc outperform GenLoc on Birt. However, when DreamLoc is executed without these additional features, its performance also drops significantly, as shown in Table LABEL:dreamloc.
Table 5. Distribution of bug reports with reproduction-related keywords
Project Total Bugs Reproduction-related Keywords Percentage (%)
@@
Table 6. Comparison of GenLoc and DreamLoc (without additional features) on the Birt Project
@@
Figure 4. Overlap Analysis between GenLoc and Non-LLM based IRBL techniques.
@@
Apart from Accuracy@k, MRR@10 and MAP@10, the number of unique bugs localized by each approach are examined. In this analysis, a bug is considered successfully localized if the correct buggy file appears within the top 10 ranked results (i.e., Accuracy@10) (Kochhar et al., 2016). Figure 4 presents the unique bugs localized by each technique under this criterion. Since GenLoc is a non-deterministic approach, its output varies across different executions due to the inherent randomness in language model sampling. To fairly evaluate GenLoc’s overall capability and compare it with existing deterministic baselines, the union of all bugs localized across three trials is considered. This union represents the total set of unique bugs that GenLoc is able to localize in at least one run, capturing the full potential of the approach rather than the outcome of a single arbitrary execution.
@@
Among all approaches, GenLoc can localize the highest number of unique bugs (601 bugs). This indicates that GenLoc not only localizes more bugs overall but also captures a unique subset of bugs missed by all other methods, highlighting its potential to complement existing methods. This is likely due to its ability to reason over natural language and semantic context, which prior techniques lack.
5.2. RQ2: Comparison with LLM-based IRBL Techniques
Table 7 reports the comparison of GenLoc against three LLM-based techniques—BRaIn, LLM-BL and Agentless—using the GHRB dataset. The results show that GenLoc consistently outperforms all baselines across every evaluation metric. It achieves the highest Accuracy@1 of 63.36%, substantially higher than other techniques, and maintains this advantage at k=5 and k=10. For ranking-based measures, GenLoc achieves a MAP@10 of 0.59 and MRR@10 of 0.69, about 20% higher than LLM-BL, while outperforming Agentless by over 75% and BRaIn by more than 100%. In addition, GenLoc localizes 5 unique bugs (Fig. 5), demonstrating that it can complement other LLM-based techniques. We observed similar results on the subset of Ye et al. dataset, with detailed results provided in the replication package.
Table 7. Performance of LLM-based IRBL Techniques